If 8,000 is invested at an annual rate of 6.5 compounded continously, find the accumulated amount after 5 1/2 years.
Answer by lwsshak3(11628)

If 8,000 is invested at an annual rate of 6.5 compounded continously, find the accumulated amount after 5 1/2 years.
**
Formula for continuous compounding: A=Pe^rt, P=initial investment, r=annual interest rate, A=amount after t years
..
A=8000*e^(.065*5.5)
A=8000*e^.3575
using calculator
A=11,438
Accumulated amount after 5 1/2 years=11,438